OpenStudy (anonymous):

If the sine of theta is (3/8) and theta is in the first quadrant, what is sin 2(theta)? a.) (3(sqrt55))/32 b.) (sqrt55)/8 c.) (2sqrt55)/32 d.) 3/8

OpenStudy (campbell_st):

1st draw a triangle and find the missing side using pythagoras  then you can use the double angle expansion of sin(2x) \[\sin(2\theta) = 2 \sin(\theta)\cos(\theta)\]

OpenStudy (anonymous):

so the answer would be c? (2sqrt55)/32?

OpenStudy (campbell_st):

its \[2 \times \frac{3}{8} \times \frac{\sqrt{55}}{8}\]